a uniform wire of resistance 6 ohms is bent in the form of a closed equilateral triangle.equivalent resistance across any two of its corners is

Grade:11

4 Answers

suvidhi mehta
6 years ago
is the answer to your question 4/3 ohms.
Khimraj
6 years ago
wire is of 6 ohm resistance then resisatnce of each side of triangle will be 2 ohms.
Then combination of resistance will be 2 ohm with parallel of series combination of 2 ohm and 2 ohm.
which is equal to 2 ohm with parralel to 4 ohm.
So equivalent resistance is 2*4/(2+4) = 4/3 ohms.
Gitanjali Rout
5 years ago
 
ire is of 6 ohm resistance then resisatnce of each side of triangle will be 2 ohms.
Then combination of resistance will be 2 ohm with parallel of series combination of 2 ohm and 2 ohm.
which is equal to 2 ohm with parralel to 4 ohm.
So equivalent resistance is 2*4/(2+4) = 4/3 ohms.
Gitanjali Rout
5 years ago
s the answer to your question 4/3 ohms.